1. You have 42.04g of a sample of carbon. Calculate the number of moles in the carbon sample
  2. 2.If you have 4.257 mol of Al(OH)3, the number of molecules you have is

3. A Canadian penny contains 0.106 mol of copper. How many atoms of copper are in a Canadian penny?

4.Iron (II) oxide is used in mascara and eye shadow because it provides a brown colour to the product.  “Pretty Eyes” mascara uses 0.15g of the Iron (II) oxide. Determine the number of “particles” (aka molecules) that would be present in the product.

5. Analysis of an ore of calcium shows that it contains 13.61g of calcium and 21.77g oxygen in a sample with a total mass of 46.28g. Calculate the mass percent of this compound.

6. Determine the % composition of Sodium Nitrate ?



1
Expert's answer
2021-04-19T04:35:54-0400

Solution:

(1):

Molar mass of carbon (C) is 12.011 g/mol.

Hence,

(42.04 g C) × (1 mol C / 12.011 g C) = 3.50 mol C

Moles of carbon = 3.50 mol


(2):

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of Al(OH)3 molecules = (4.257 mol) × (6.022×1023 molecules / 1 mol) = 2.56×1024 molecules

Number of Al(OH)3 molecules = 2.56×1024 molecules


(3):

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of copper atoms = (0.106 mol) × (6.022×1023 atoms / 1 mol) = 6.38×1022 atoms

Number of copper atoms = 6.38×1022 atoms


(4):

Iron (II) oxide - FeO

The molar mass of FeO is 71.844 g/mol.

Hence,

(0.15 g FeO) × (1 mol FeO / 71.844 g FeO) = 0.002088 mol

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of FeO molecues = (0.002088 mol) × (6.022×1023 molecules / 1 mol) = 1.257×1021 molecules

Number of FeO molecues = 1.257×1021 molecules


(5):

%Ca = (Mass of Ca / Total mass of sample) × 100%

%Ca = (13.61 g / 46.28 g) × 100% = 29.41%

%Ca = 29.41%

%O = (Mass of O / Total mass of sample) × 100%

%O = (21.77 g / 46.28 g) × 100% = 47.04%

%O = 47.04%


(6):

Sodium nitrate - NaNO3

The molar mass of NaNO3 is 84.995 g/mol.

The molar mass of Na is 22.9898 g/mol.

The molar mass of N is 14.0067 g/mol.

The molar mass of O is 15.999 g/mol.

Hence,

%Na = (22.9898 / 84.995) × 100% = 27.05%

%N = (14.0067 / 84.995) × 100% = 16.48%

%O = (3×15.999 / 84.995) × 100% = 56.47%

%Na = 27.05%

%N = 16.48%

%O = 56.47%


Answers:

1) Moles of carbon = 3.50 mol;

2) Number of Al(OH)3 molecules = 2.56×1024 molecules;

3) Number of copper atoms = 6.38×1022 atoms;

4) Number of FeO molecues = 1.257×1021 molecules;

5) %Ca = 29.41%; %O = 47.04%;

6) %Na = 27.05%; %N = 16.48%; %O = 56.47%.